The difference of squares of two numbers is 180. The square of the smaller number is 8 times the larger number. Find the two numbers.
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us say, the larger and smaller number bexandyrespectively. As per the question given,
x^2-y^2= 180
andy^2= 8x
⇒x^2- 8x= 180

⇒x^2- 8x- 180 = 0
⇒x^2- 18x+ 10x- 180 = 0
⇒x(x- 18) +10(x- 18) = 0
⇒(x- 18)(x+ 10) = 0
⇒x= 18, -10

However, the larger number cannot considered as negative number, as 8 times of the larger number will be negative and hence, the square of the smaller number will be negative which is not possible.
Therefore, the larger number will be 18 only.
x= 18

∴y^2= 8x = 8 × 18 = 144
⇒y= ±√144 = ±12
∴Smaller number = ±12
Therefore, the numbers are 18 and 12 or 18 and -12.

Thanks